# “Quantum Geometric Nesting” and Solvable Model Flat-Band Systems

### v2

@@ V.3 Non-symmetry-related cases: An engineered model for pair-density-wave
with $\hat{\bar{c}}^{\dagger}_{\bm{k}+\bm{Q},\nu}=\sum_{\nu}\hat{\bar{c}}^{\dagger}_{\bm{k}+\bm{Q},\mu}P_{\mu\nu}(\bm{k})$ the projection of the orbital operators onto the flat bands (see Eq. (9)). Projecting $\hat{H}_{\text{int}}$ in Eq. (31) to the flat band Hilbert space results in trivial one-body (Hartree-Fock-type) terms that act as a chemical potential on the flat bands. The triviality of the one-body terms generalizes the “uniform pairing condition” Tovmasyan et al. (2016); Herzog-Arbeitman et al. (2022) utilized in attractive Hubbard models (see Sec. II D of the SM).
Adopting the methods in Sec. IV, we can exactly solve the few-body excitations in an artificial setup by setting the $W\to 0$ and $\Delta\to\infty$ with the quantum geometry $U_{\mu n}(\bm{k})$ remaining fixed. First, we consider the charge ±$\pm 1$ fermionic quasi-particles above the ground states. We find that these excitations are spin-degenerate, gapped, and flat, with energy $E^{(\pm 1)}(\mathbf{p})=3V$. Next, we consider a Cooper pair excitation. This spectrum is also analytically solvable since the two-particle scattering matrix can be diagonalized by generalizing the approach of Ref. Herzog-Arbeitman et al. (2022) to extract a low-rank Hamiltonian for the pair binding energies. We find that there are two bound modes - the Goldstone mode of the PDW condensate, and a higher-energy Leggett mode - below the flat and macroscopically degenerate two-particle continuum at $6V$. The Goldstone mode has dispersion
@@ VI Possible Extensions
which realizes an obstructed atomic limit Song et al. (2020) with lower-bounded quantum geometry Herzog-Arbeitman et al. (2022); Espinosa-Champo and Naumis (2023). As discussed in the previous section, the time-reversal symmetry and spin conservation imply perfect QGN in the p-p channel, suggesting a singlet $s$-wave superconducting groundstate Herzog-Arbeitman et al. (2022). However, we find that this result can be promoted to a singlet $p$-wave OP
@@
Acknowledgement: We would like to thank Erez Berg, Daniel Parker, Junkai Dong, Ashvin Vishwanath, Chaitanya Murthy, Biao Lian, Ben Feldman, Oskar Vafek, Jiabin Yu, Zhi-da Song, and Päivi Törmä for helpful discussions. We also thank anonymous reviewers for inspiring questions and comments. ZH was funded, in part, by a QuantEmX grant from ICAM and the Gordon and Betty Moore Foundation through Grant GBMF9616. BAB was primarily supported by the the Simons Investigator Grant No. 404513. JHA was supported by a Hertz Fellowship. BAB and JHA received additional support from the Gordon and Betty Moore Foundation through Grant No. GBMF8685 towards the Princeton theory program, Office of Naval Research (ONR Grant No. N00014-20-1-2303), BSF Israel US foundation No. 2018226 and NSF-MERSEC DMR-2011750, Princeton Global Scholar and the European Union’s Horizon 2020 research and innovation program under Grant Agreement No 101017733 and from the European Research Council (ERC). SAK and ZH were supported, in part, by the Department of Energy, Office of Basic Energy Sciences, under contract DE-AC0276SF00515 at Stanford.

### v3

@@ V.3 Non-symmetry-related cases: An engineered model for pair-density-wave
with $\hat{\bar{c}}^{\dagger}_{\bm{k}+\bm{Q},\nu}=\sum_{\nu}\hat{\bar{c}}^{\dagger}_{\bm{k}+\bm{Q},\mu}P_{\mu\nu}(\bm{k})$ the projection of the orbital operators onto the flat bands (see Eq. (9)). Projecting $\hat{H}_{\text{int}}$ in Eq. (31) to the flat band Hilbert space results in trivial one-body (Hartree-Fock-type) terms that act as a chemical potential on the flat bands. The triviality of the one-body terms generalizes the “uniform pairing condition” Tovmasyan et al. (2016); Herzog-Arbeitman et al. (2022a) utilized in attractive Hubbard models (see Sec. II D of the SM).
Adopting the methods in Sec. IV, we can exactly solve the few-body excitations in an artificial setup by setting the $W\to 0$ and $\Delta\to\infty$ with the quantum geometry $U_{\mu n}(\bm{k})$ remaining fixed. First, we consider the charge ±$\pm 1$ fermionic quasi-particles above the ground states. We find that these excitations are spin-degenerate, gapped, and flat, with energy $E^{(\pm 1)}(\mathbf{p})=3V$. Next, we consider a Cooper pair excitation. This spectrum is also analytically solvable since the two-particle scattering matrix can be diagonalized by generalizing the approach of Ref. Herzog-Arbeitman et al. (2022a) to extract a low-rank Hamiltonian for the pair binding energies. We find that there are two bound modes - the Goldstone mode of the PDW condensate, and a higher-energy Leggett mode - below the flat and macroscopically degenerate two-particle continuum at $6V$. The Goldstone mode has dispersion
@@ VI Possible Extensions
which realizes an obstructed atomic limit Song et al. (2020) with lower-bounded quantum geometry Herzog-Arbeitman et al. (2022b); de Jesús Espinosa-Champo and Naumis (2023). As discussed in the previous section, the time-reversal symmetry and spin conservation imply perfect QGN in the p-p channel, suggesting a singlet $s$-wave superconducting groundstate Herzog-Arbeitman et al. (2022a). However, we find that this result can be promoted to a singlet $p$-wave OP
@@
Acknowledgement: We would like to thank Erez Berg, Daniel Parker, Junkai Dong, Ashvin Vishwanath, Chaitanya Murthy, Biao Lian, Ben Feldman, Oskar Vafek, Jiabin Yu, Zhi-Da Song, and Päivi Törmä for helpful discussions. ZH was funded, in part, by a QuantEmX grant from ICAM and the Gordon and Betty Moore Foundation through Grant GBMF9616. BAB was primarily supported by the the Simons Investigator Grant No. 404513. JHA was supported by a Hertz Fellowship. BAB and JHA received additional support from the Gordon and Betty Moore Foundation through Grant No. GBMF8685 towards the Princeton theory program, Office of Naval Research (ONR Grant No. N00014-20-1-2303), BSF Israel US foundation No. 2018226 and NSF-MERSEC DMR-2011750, Princeton Global Scholar and the European Union’s Horizon 2020 research and innovation program under Grant Agreement No 101017733 and from the European Research Council (ERC). SAK and ZH were supported, in part, by the Department of Energy, Office of Basic Energy Sciences, under contract DE-AC0276SF00515 at Stanford.